OpenStudy (anonymous):

Does anyone know how to solve c=1.25+0.30x c=1.40+0.25x?

OpenStudy (anonymous):

Sub the equations into each other. Sub the first c into the second one to get: 1.25+0.3x = 1.4+0.25x 0.05x = 0.15 x = 3

OpenStudy (anonymous):

oh ok, thank-you Riddellikins

OpenStudy (anonymous):

You welcome